# Theory of Mind abilities of Large Language Models in Human-Robot Interaction : An Illusion?

### 10 Jan 2024 v1

@@ C.1. Fetch Robot Domain
Description: Fetch is a robot that can carry objects (pick / place) and move from one location to another. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Goal: The robot has to take the block b1 to location loc2. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot tucks its arms and lowers itself. It then picks up the block b1. Finally, the robot moves to location loc2. Question 1 : Imagine you are the human observer, would you find such a plan explicable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) The plan is not explicable because the robot tucked its arm before moving, which seems unnecessary. 2) The plan is not explicable because the robot picked up block b1 which seems unnecessary. 3) The plan is not explicable because no plan is explicable.
@@
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. Question 1 : Would you find such a partial plan legible? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) The plan is legible because the robot made its goal of taking the block to location loc2 more legible. 2) The plan is legible because the robot made its goal of picking up the block more legible. 3) The plan is legible because all robot plans are legible.
@@
Description: Fetch is a robot that can carry objects from one location to another. Initial state: There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location locX can be reached from location loc1 using two paths, one by taking five steps to left and the other one by taking five steps to right. Goal: The robot has to pick the block b1 and take it to location locX. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step towards left. Question 1 : Would you find such a partial plan predictable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) because the robot will go to locX from left path. 2) because the robot will go to locX from right path. 3) because all paths are predictable.
@@
Description: Fetch is a robot that can carry objects from one location to another. Fetch robot’s design requires it to tuck its arms and lower its torso or crouch before moving. There are three locations: loc1, loc2, and loc3. Initial state: There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the block b1. At loc1 is the agent is at a distance of 10 steps from loc2 and loc3. It takes 3 steps forward and is 7 steps away from loc2 as well as loc3. Question 1 : Would you find such a partial plan obfuscatory? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) because the robot can possibly go to either location loc2 or location loc3. 2) because the robot will only go to location loc2. 3) because all plans are obfuscatory.
@@ C.2. Passage Gridworld Domain
Description: Consider a 4x4 square grid with each cell numbered as (row, column), the robot needs to travel from top left cell 1 (1,1) to its goal at bottom right cell 15 (4,3). The human observer (you) expects the robot to take the shortest path by going DOWN 3 steps to row 4 (reach 4, 1), and then RIGHT 2 steps to column 3 (reach 4,3). Constraint: There are blockades in columns 1, 2 and 3, that the human observer (you) do not know of, but the robot does. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan: The robot goes RIGHT 3 steps in row 1 (reach 1,4), goes DOWN 3 steps in column 4 (reach 4,4), and LEFT 1 step in row 4 (4,3). Question 1: Imagine you are the human observer, would you find such a plan explicable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human does not know that the path from row 1 to row 4 in column 1 is blocked. 2) Because the robot does not know that the path from row 1 to row 4 in column 1 is blocked. 3) Because the human knows that the path from row 1 to row 4 in column 1 is blocked.
@@
Description: Consider a 4x4 square grid with each cell numbered as (row, column), and the robot starts to travel from top left cell 1 (1,1). Constraint: There are blockades in columns 1, 2 and 3, that the human observer (you) do not know of, but the robot does. Goals: The robot has to reach either cell 12 (3,4) or cell 16 (4,4) Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it goes DOWN 3 steps to reach cell (4, 1). Question 1 : Imagine you are the human observer in this case. Would you find such a plan legible? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human will know that the agent is going to go RIGHT to cell 16 (4,4). 2) Because the human will know that the agent is going to RIGHT and then UP to cell 12 (4,3). 3) Because all plans are equally legible.
@@
Description: Consider a 4x4 square grid with each cell numbered as (row, column), and the robot starts to travel from top left cell 1 (1,1). Constraint: There are blockades in columns 1, 2 and 4, that both, the human observer (you) and the robot know of. Goals: The robot has to reach cell 15 (4,3). Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it goes RIGHT 3 steps to reach (1,3). Question 1: Imagine you are a human in this case. Would you find such a partial plan predictable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot will now go DOWN in column 3 to reach the goal. 2) Because the robot should have first gone DOWN in column 1 to reach row 4. 3) Because all plans are equally predictable.
@@
Description: Consider a 4x4 square grid with each cell numbered as (row, column), and the robot starts to travel from top left cell 1 (1,1). Constraint: There are blockades in columns 1, 2 and 3, that both, the human observer (you) and the robot know of. Goals: The robot has to reach either cell 12 (3,4) or cell 16 (4,4) Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : In the robot’s partial plan, it goes RIGHT 3 steps to reach cell (1,4). Question 1: Imagine you are a human in this case. Would you find such a plan obfuscatory? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot can reach any of the two goal cells (4,3) and (4,4). 2) Because the robot can only reach cell (4,4). 3) Because the robot does not know of the blockades and chose the wrong plan.
@@ C.3. Environment Design Domain
Description: There is a 3x3 square grid numbered as (row, column) = (1,1) the bottom left cell and (3,3) is the top right cell. The robot needs to travel from cell 1 (1,1) to achieve two goals G1, placed at (3,1) and G2, placed at (3,2). The robot cannot go through cells that have an obstacle. The robot can go UP, DOWN, LEFT or RIGHT. Finally, there may be objects placed in one or more cells, and the agent will incur a very high cost on visiting these two cells. Suppose there are two different instantiations of this grid based on how these obstacles are placed in the environment: Setup A : No obstacles. Setup B : Obstacle T1 at (2,1) and T2 at (2,2). Goals : The robot needs to find a plan to reach both goals, G1 and G2 (in any order). Definition: An explicable environment setup is one where the robot can take a single plan that is easily expected by a human observer (you). If you look at the environment and can easily identify a plan for the robot to follow, then the environment setup is explicable. Question 1 : For Setup A there are multiple valid plans possible, for example to goto cells (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Another shorter plan can be (1,1) UP (2,1) UP (3,1) RIGHT (3,2). For Setup B, there is only one valid plan i.e. (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Given the above description of the two setups, which environment do you believe is designed to be explicable? Give your answer only as Setup A or Setup B. ——————- Question 2 : Why do you think Setup A is not explicable?Choose one of the following reasons for your answer. Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because I believe the agent should take the shorter plan which is costlier in Setup A. 2) Because I believe the agent should take the longer plan which is cheaper in Setup A. 3) Because such an environment does not exist.
@@
Description: There is a 3x3 square grid numbered as (row, column) = (1,1) the bottom left cell and (3,3) is the top right cell. The robot needs to travel from cell 1 (1,1) to achieve two goals G1, placed at (3,1) and G2, placed at (3,2). The robot cannot go through cells that have an obstacle. The robot can go UP, DOWN, LEFT or RIGHT. Finally, there may be objects placed in one or more cells, and the agent will incur a very high cost on visiting these two cells. Suppose there are two different instantiations of this grid based on how these obstacles are placed in the environment: Setup A : No obstacles. Setup B : Obstacle T1 at (2,1) and T2 at (2,2). Goals : The robot needs to find a plan to reach both goals, G1 and G2 (in any order). Definition: A legible environment setup is one where the number of plans a robot can take to achieve the goals is the minimum. Question 1: For Setup A there are multiple plans possible, for example to goto cells (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Another shorter plan can be (1,1) UP (2,1) UP (3,1) RIGHT (3,2). For Setup B, there is only one valid plan i.e. (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Given the above description of the two setups, which environment do you believe is designed to be legible? Give your answer only as Setup A or Setup B. ——————- Question 2 : Why do you think Setup A is not legible? Choose one of the following reasons for your answer. Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because there are multiple plans that achieve the goal in Setup A. 2) Because there is a single plan that achieves the goal in Setup A. 3) Because such an environment does not exist.
@@
Description: There is a 3x3 square grid numbered as (row, column) = (1,1) the bottom left cell and (3,3) is the top right cell. The robot needs to travel from cell 1 (1,1) to achieve two goals G1, placed at (3,1) and G2, placed at (3,2). The robot cannot go through cells that have an obstacle. The robot can go UP, DOWN, LEFT or RIGHT. Finally, there may be objects placed in one or more cells, and the agent will incur a very high cost on visiting these two cells. Suppose there are two different instantiations of this grid based on how these obstacles are placed in the environment: Setup A : No obstacles. Setup B : Obstacle T1 at (2,1) and T2 at (2,2). Goals : The robot needs to find a plan to reach both goals, G1 and G2 (in any order). Definition: A predictable environment setup is one where the set of possible plans for the robot is as low as possible. Question 1: For Setup A there are multiple plans possible, for example to goto cells (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Another shorter plan can be (1,1) UP (2,1) UP (3,1) RIGHT (3,2). For Setup B, there is only one valid plan i.e. (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Given the above description of the two setups, which environment do you believe is designed to be predictable? Give your answer only as Setup A or Setup B. ——————- Question 2 : Why do you think Setup A is not predictable? Choose one of the following reasons for your answer. Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because there are multiple plans that can be executed in Setup A, unlike setup B. 2) Because there is a single plan that can be extracted in Setup A, unlike Setup B. 3) Because such an environment does not exist.
@@
Description: There is a 3x3 square grid numbered as (row, column) = (1,1) the bottom left cell and (3,3) is the top right cell. The robot needs to travel from cell 1 (1,1) to achieve two goals G1, placed at (3,1) and G2, placed at (3,2). The robot cannot go through cells that have an obstacle. The robot can go UP, DOWN, LEFT or RIGHT. Finally, there may be objects placed in one or more cells, and the agent will incur a very high cost on visiting these two cells. Suppose there is only one instantiation of this grid based on how these obstacles are placed in the environment: Setup A : No obstacles. Goals : The robot needs to find a plan to reach one of the two goals, G1 and G2. Definition: An environment is designed for obfuscation when all the plan completions are equally worse for all the agent goals. This is useful when the agent wants to achieve a certain goal say G1 but does not want the observer (you) to realize which among set of possible goals it wants to achieve. An environment designed for obfuscations allows for plans that lets the agent hide which goal it wants to achieve for as long as possible. Question 1 : For the setup there are multiple plans possible, for example to goto cells (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Another shorter plan can be (1,1) UP (2,1) UP (3,1) RIGHT (3,2). Do you think that the environment is designed for obfuscation? Give your answer only as Setup A or Setup B. ——————- Question 2 : Why do you think the setup is not designed to be obfuscatory? Choose one of the following reasons for your answer. Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because there are multiple plans that can quickly reveal the agent’s goal 2) Because there is only one plan that can quickly reveal the agent’s goal. 3) Because such an environment does not exist.

### 17 Jan 2024 v2

@@ C.1. Fetch Robot Domain
Description: Fetch is a robot that can carry objects (pick / place) and move from one location to another. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Goal: The robot has to take the block b1 to location loc2. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot tucks its arms and lowers itself. It then picks up the block b1. Finally, the robot moves to location loc2. The robot has a label saying ’Explicable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a plan explicable? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) The plan is not explicable because the robot tucked its arm before moving, which seems unnecessary. 2) The plan is not explicable because the robot picked up block b1 which seems unnecessary. 3) The plan is not explicable because no plan is explicable.
@@
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. The robot has a label saying ’Not a Legible Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) The plan is legible because the robot made its goal of taking the block to location loc2 more legible. 2) The plan is legible because the robot made its goal of picking up the block more legible. 3) The plan is legible because all robot plans are legible.
@@
Description: Fetch is a robot that can carry objects from one location to another. Initial state: There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location locX can be reached from location loc1 using two paths, one by taking five steps to left and the other one by taking five steps to right. Goal: The robot has to pick the block b1 and take it to location locX. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step towards left. The robot has a label saying ’Not a Predictable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan predictable? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) because the robot will go to locX from left path. 2) because the robot will go to locX from right path. 3) because all paths are predictable.
@@
Description: Fetch is a robot that can carry objects from one location to another. Fetch robot’s design requires it to tuck its arms and lower its torso or crouch before moving. There are three locations: loc1, loc2, and loc3. Initial state: There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the block b1. At loc1 is the agent is at a distance of 10 steps from loc2 and loc3. It takes 3 steps forward and is 7 steps away from loc2 as well as loc3. The robot has a label saying ’Not an Obfuscatory Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan obfuscatory? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) because the robot can possibly go to either location loc2 or location loc3. 2) because the robot will only go to location loc2. 3) because all plans are obfuscatory.
@@ C.2. Passage Gridworld Domain
Description: Consider a 4x4 square grid with each cell numbered as (row, column), the robot needs to travel from top left cell 1 (1,1) to its goal at bottom right cell 15 (4,3). The human observer (you) expects the robot to take the shortest path by going DOWN 3 steps to row 4 (reach 4, 1), and then RIGHT 2 steps to column 3 (reach 4,3). Constraint: There are blockades in columns 1, 2 and 3, that the human observer (you) do not know of, but the robot does. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan: The robot goes RIGHT 3 steps in row 1 (reach 1,4), goes DOWN 3 steps in column 4 (reach 4,4), and LEFT 1 step in row 4 (4,3). The robot has a label saying ’Explicable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1: Imagine you are the human observer. Would you find such a plan explicable? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human does not know that the path from row 1 to row 4 in column 1 is blocked. 2) Because the robot does not know that the path from row 1 to row 4 in column 1 is blocked. 3) Because the human knows that the path from row 1 to row 4 in column 1 is blocked.
@@
Description: Consider a 4x4 square grid with each cell numbered as (row, column), and the robot starts to travel from top left cell 1 (1,1). Constraint: There are blockades in columns 1, 2 and 3, that the human observer (you) do not know of, but the robot does. Goals: The robot has to reach either cell 12 (3,4) or cell 16 (4,4) Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it goes DOWN 3 steps to reach cell (4, 1). The robot has a label saying ’Not a Legible Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a plan legible? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human will know that the agent is going to go RIGHT to cell 16 (4,4). 2) Because the human will know that the agent is going to RIGHT and then UP to cell 12 (4,3). 3) Because all plans are equally legible.
@@
Description: Consider a 4x4 square grid with each cell numbered as (row, column), and the robot starts to travel from top left cell 1 (1,1). Constraint: There are blockades in columns 1, 2 and 4, that both, the human observer (you) and the robot know of. Goals: The robot has to reach cell 15 (4,3). Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it goes RIGHT 3 steps to reach (1,3). The robot has a label saying ’Not a Predictable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1: Imagine you are the human observer. Would you find such a partial plan predictable? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot will now go DOWN in column 3 to reach the goal. 2) Because the robot should have first gone DOWN in column 1 to reach row 4. 3) Because all plans are equally predictable.
@@
Description: Consider a 4x4 square grid with each cell numbered as (row, column), and the robot starts to travel from top left cell 1 (1,1). Constraint: There are blockades in columns 1, 2 and 3, that both, the human observer (you) and the robot know of. Goals: The robot has to reach either cell 12 (3,4) or cell 16 (4,4) Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : In the robot’s partial plan, it goes RIGHT 3 steps to reach cell (1,4). The robot has a label saying ’Not an Obfuscatory Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1:Imagine you are the human observer. Would you find such a plan obfuscatory? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot can reach any of the two goal cells (4,3) and (4,4). 2) Because the robot can only reach cell (4,4). 3) Because the robot does not know of the blockades and chose the wrong plan.
@@ C.3. Environment Design Domain
Description: There is a 3x3 square grid numbered as (row, column) = (1,1) the bottom left cell and (3,3) is the top right cell. The robot needs to travel from cell 1 (1,1) to achieve two goals G1, placed at (3,1) and G2, placed at (3,2). The robot cannot go through cells that have an obstacle. The robot can go UP, DOWN, LEFT or RIGHT. Finally, there may be objects placed in one or more cells, and the agent will incur a very high cost on visiting these two cells. Suppose there are two different instantiations of this grid based on how these obstacles are placed in the environment: Setup A : No obstacles. Setup B : Obstacle T1 at (2,1) and T2 at (2,2). Goals : The robot needs to find a plan to reach both goals, G1 and G2 (in any order). Definition: An explicable environment setup is one where the robot can take a single plan that is easily expected by a human observer (you). If you look at the environment and can easily identify a plan for the robot to follow, then the environment setup is explicable. Question 1 : For Setup A there are multiple valid plans possible, for example to goto cells (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Another shorter plan can be (1,1) UP (2,1) UP (3,1) RIGHT (3,2). For Setup B, there is only one valid plan i.e. (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Setup B has a label saying ’Inexplicable Setup’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the setup before. Imagine you are the human observer. Given the above description of the two setups, which environment do you believe is designed to be explicable? Give your answer as ‘Setup A’ or ‘Setup B’ only. ——————- Question 2 : Why do you think Setup A is not explicable?Choose one of the following reasons for your answer. Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because I believe the agent should take the shorter plan which is costlier in Setup A. 2) Because I believe the agent should take the longer plan which is cheaper in Setup A. 3) Because such an environment does not exist.
@@
Description: There is a 3x3 square grid numbered as (row, column) = (1,1) the bottom left cell and (3,3) is the top right cell. The robot needs to travel from cell 1 (1,1) to achieve two goals G1, placed at (3,1) and G2, placed at (3,2). The robot cannot go through cells that have an obstacle. The robot can go UP, DOWN, LEFT or RIGHT. Finally, there may be objects placed in one or more cells, and the agent will incur a very high cost on visiting these two cells. Suppose there are two different instantiations of this grid based on how these obstacles are placed in the environment: Setup A : No obstacles. Setup B : Obstacle T1 at (2,1) and T2 at (2,2). Goals : The robot needs to find a plan to reach both goals, G1 and G2 (in any order). Definition: A legible environment setup is one where the number of plans a robot can take to achieve the goals is the minimum. Question 1: For Setup A there are multiple plans possible, for example to goto cells (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Another shorter plan can be (1,1) UP (2,1) UP (3,1) RIGHT (3,2). For Setup B, there is only one valid plan i.e. (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Setup B has a label saying ’Not a Legible Setup’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the setup before. Imagine you are the human observer. Given the above description of the two setups, which environment do you believe is designed to be legible? Give your answer as ‘Setup A’ or ‘Setup B’ only. ——————- Question 2 : Why do you think Setup A is not legible? Choose one of the following reasons for your answer. Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because there are multiple plans that achieve the goal in Setup A. 2) Because there is a single plan that achieves the goal in Setup A. 3) Because such an environment does not exist.
@@
Description: There is a 3x3 square grid numbered as (row, column) = (1,1) the bottom left cell and (3,3) is the top right cell. The robot needs to travel from cell 1 (1,1) to achieve two goals G1, placed at (3,1) and G2, placed at (3,2). The robot cannot go through cells that have an obstacle. The robot can go UP, DOWN, LEFT or RIGHT. Finally, there may be objects placed in one or more cells, and the agent will incur a very high cost on visiting these two cells. Suppose there are two different instantiations of this grid based on how these obstacles are placed in the environment: Setup A : No obstacles. Setup B : Obstacle T1 at (2,1) and T2 at (2,2). Goals : The robot needs to find a plan to reach both goals, G1 and G2 (in any order). Definition: A predictable environment setup is one where the set of possible plans for the robot is as low as possible. Setup B has a label saying ’Not a Predictable Setup’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the setup before. Question 1: For Setup A there are multiple plans possible, for example to goto cells (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Another shorter plan can be (1,1) UP (2,1) UP (3,1) RIGHT (3,2). For Setup B, there is only one valid plan i.e. (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Imagine you are the human observer. Given the above description of the two setups, which environment do you believe is designed to be predictable? Give your answer as ‘Setup A’ or ‘Setup B’ only. ——————- Question 2 : Why do you think Setup A is not predictable? Choose one of the following reasons for your answer. Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because there are multiple plans that can be executed in Setup A, unlike setup B. 2) Because there is a single plan that can be extracted in Setup A, unlike Setup B. 3) Because such an environment does not exist.
@@
Description: There is a 3x3 square grid numbered as (row, column) = (1,1) the bottom left cell and (3,3) is the top right cell. The robot needs to travel from cell 1 (1,1) to achieve two goals G1, placed at (3,1) and G2, placed at (3,2). The robot cannot go through cells that have an obstacle. The robot can go UP, DOWN, LEFT or RIGHT. Finally, there may be objects placed in one or more cells, and the agent will incur a very high cost on visiting these two cells. Suppose there is only one instantiation of this grid based on how these obstacles are placed in the environment: Setup A : No obstacles. Goals : The robot needs to find a plan to reach one of the two goals, G1 and G2. Definition: An environment is designed for obfuscation when all the plan completions are equally worse for all the agent goals. This is useful when the agent wants to achieve a certain goal say G1 but does not want the observer (you) to realize which among set of possible goals it wants to achieve. An environment designed for obfuscations allows for plans that lets the agent hide which goal it wants to achieve for as long as possible. Setup A has a label saying ’Not an Obfuscatory Setup’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the setup before. Question 1 : For the setup there are multiple plans possible, for example to goto cells (1,1) RIGHT (1, 2) RIGHT (1,3) UP (2, 3), LEFT(2, 2), LEFT (2, 1) UP (3, 1) RIGHT (3, 2). Another shorter plan can be (1,1) UP (2,1) UP (3,1) RIGHT (3,2). Imagine you are the human observer. Do you think that the environment is designed for obfuscation? Give your answer as ‘Setup A’ or ‘Setup B’ only. ——————- Question 2 : Why do you think the setup is not designed to be obfuscatory? Choose one of the following reasons for your answer. Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because there are multiple plans that can quickly reveal the agent’s goal 2) Because there is only one plan that can quickly reveal the agent’s goal. 3) Because such an environment does not exist.